# F&O Expiry vs. First-Day SIPs: A 22-Year Analysis of Timing Advantages in India’s Nifty 50

### v1

@@ II Literature Review
    Timing of SIP Contributions within Monthly Cycles: Most studies focus on the benefits of SIPs over lump-sum investments but overlook intra-month contribution timing (e.g., beginning of the month vs. expiry days), especially in the context of Indian market microstructures.
@@
    Structural and Institutional Influences: The potential distortions caused by concentrated retail flows on specific trading days, and their interaction with institutional behavior, remain underexplored in long-term wealth accumulation studies.
These gaps highlight the need for a more nuanced and data-driven understanding of SIP timing, and market frictions-specific behavior—particularly in emerging markets like India, where retail participation is growing rapidly.
III Methodology
This study evaluates the impact of timing and duration on the performance of Systematic Investment Plans (SIPs) using historical data from the Nifty 50 index over the 2003–2024 period extracted from NSE India [20]. This window is selected to ensure data completeness following the introduction of stock and index derivatives (Futures & Options) in India during 2000–2001[22]. While direct investment in the Nifty 50 is not feasible, it serves as a representative benchmark for equity market performance and is closely tracked by index funds and ETFs [1]. To ensure realistic and replicable calculations, daily closing prices are used for all return computations, as intraday highs and lows are not consistently accessible and are operationally impractical for SIP execution—even for high-frequency systems [33]. All SIPs are simulated using calendar-year windows (January to December), maintaining uniformity across scenarios. A detailed mapping of monthly first trading days and F&O expiry dates used for investment timing is provided in Table I.
@@ III Methodology
TABLE I: First trading date and F&O expiry date (in parentheses) for each month
III-A Investment Durations
III-B Investment Timing
III-C Return Computation and Statistical Evaluation
  • $I^{E}_{i}$: Closing index value on the F&O expiry day of the $i^{\text{th}}$ month.
  • $q^{k}_{i}=\frac{m}{I^{k}_{i}}$: Units purchased in month $i$ under strategy $k$.
$Q^{k}=\sum_{i=1}^{12N}q^{k}_{i}$
The final value of the investment at the end of the horizon is:
$F^{k}=I^{L}_{12N}\cdot Q^{k},$ (1)
where $I^{L}_{12N}$ denotes the index closing value on the last trading day of December in the terminal year. Specific maturity dates are listed in Table II.
TABLE II: Last trading date of year (2003-2024)
For each of the 36 investment plans across all durations and schedules (FTD and EXP), we compute two separate CAGR values—CAGR_f and CAGR_e—to compare the impact of timing on final returns. This allows us to examine whether entry-day strategy has a systematic influence on performance, particularly in the presence of market volatility and cyclicality. To assess the statistical significance of observed differences, we conduct two-sample paired t-tests for each investment horizon (1-year, 3-year, 5-year, and 10-year periods) using Excel’s built-in T.TEST function. Specifically, the test is performed using the syntax T.TEST(CAGR_e, CAGR_f, 2, 1), which evaluates whether the mean difference between expiry-day (EXP) and first-day (FTD) CAGR distributions is statistically significant under a two-tailed assumption with paired data.
IV Results and Analysis
IV-A Theoretical
Lemma 1.
(CAGR Dependency on Duration and Index Performance) The Compound Annual Growth Rate (CAGR) return of a Nifty 50 index fund does not depend on the invested amount but is inherently dependent on the investment duration and the performance of the underlying index over that period.
Proof.
Put $q^{k}_{i}=\frac{m}{I^{k}_{i}}$ in (2) gives
IV-B Empirical Result
This subsection synthesizes the empirical findings from multiple time horizons—ranging from 1-year to 20-year investment windows—based on SIP execution strategies. The data exhibits nuanced but consistent trends.
IV-B1 One-Year Investment Returns
This subsection presents a year-by-year analysis of one-year Systematic Investment Plan (SIP) returns for the Nifty 50 index between 2003 and 2024. Each year’s investment spans from January to December, using earlier mentioend two strategies. The Compounded Annual Growth Rate (CAGR) serves as the return metric, and the paired two-sample t-test is used to evaluate the statistical significance of differences between the two strategies on an annual basis.
The data exhibit considerable inter-annual variability, reflecting the inherent volatility in equity markets. Years such as 2003, 2005, 2006, 2007, 2009 & 2020, recorded CAGRs exceeding 20% under both strategies. In 2003, the first-day investment yielded a CAGR of 60.43% compared to 62.94% from expiry-based SIPs. See Table III and Figure 1 for details. Conversely, bearish years such as 2008, 2011 and 2015 experienced negative returns. The 2008 financial crisis led to nearly identical CAGRs: -30.08% for the first-day strategy and -29.99% for expiry-day SIPs, with an insignificant difference. In 2011, both strategies again posted losses, with expiry-day investments slightly outperforming (-12.58% vs. -14.81%), though this difference was not statistically significant. Same has repeated in 2015 with smaller difference in CAGR. Over the entire 22-year span, the average absolute CAGR difference between the two strategies remained modest—typically within a $\pm$2range. Positive CAGR differences in favor of expiry-based investing were observed more than 80% of the years. In 2004, 2005, and 2022, expiry-based SIPs outperformed by more than 1.55%. Years like 2003, 2011, and 2013 showed CAGR differences exceeding 2%. Meanwhile, in stable years such as 2015 and 2019, both strategies produced nearly identical outcomes, emphasizing the limited impact of timing in less volatile periods.
The t-test for this one year horizon returned a p-value of 0.0036, indicating a statistically significant difference at the 1% level. This suggests that in years of strong market rebounds, even minor timing adjustments can have measurable effects. The analysis reveals that even within a relatively short one-year investment horizon, SIPs timed around the F&O expiry date consistently delivered marginally superior returns compared to first-day-of-month investments. This suggests that expiry-based timing is not just a random anomaly but may be capturing systematic behavioral or structural market effects. While the average annual CAGR difference remained within $\pm$2%, the persistent directional advantage in favor of expiry-based SIPs highlights its potential as a tactical enhancement, especially in turbulent or momentum-driven markets. Thus, even for short-term SIPs, aligning investments with expiry dates may offer a subtle yet repeatable edge — reinforcing the case for expiry-aware SIP execution strategies.
Figure 1: Year-wise difference in one-year SIP returns between EXP-SIP and FTD-SIP
From Year To Year Years CAGR (F) CAGR (E) Difference t test
@@ IV-B1 One-Year Investment Returns
2005 2014 10 6.76 6.83 0.07
2015 2024 10 6.99 7.04 0.05 0.0687 $>$ 0.05
2005 2024 20 6.65 6.68 0.03
TABLE III: Multi-Year SIP Investment Performance: First Day vs F&O Expiry
IV-B2 Three-Year Investment Returns
The analysis of three-year rolling SIP investments reveals a clearer picture of performance differentials between the two strategies by reducing the noise of short-term market volatility. Over the seven three-year periods from 2004–2006 to 2022–2024, both the first-day and expiry-day strategies showed reasonably strong growth, with CAGR values ranging from approximately 3% to 21%. Notably, the expiry-based strategy outperformed the first-day strategy in six out of seven periods, with CAGR advantages typically ranging between 0.04% and 0.63%, see Figure 2 and Table III for details. While these differences may appear modest in absolute terms, their consistency over multiple periods warrants attention. Statistical testing via two-sample t-tests for this investment horizon adds further depth to the interpretation.The difference in returns was not only higher but also statistically significant, with p-values ($=0.029$) below the conventional 0.05 threshold.The p-value confirms that this difference is unlikely to have occurred by chance. Importantly, unlike the more erratic one-year results, the three-year intervals display a smoother, more predictable trend, reinforcing the importance of extending investment horizons in SIP-based strategies. The observed edge in expiry-day investments during multiple periods may reflect favorable end-of-month liquidity patterns or systematic institutional behaviors around derivatives expiry that influence price levels. The three-year investment analysis strengthens the case for expiry-based SIP timing. The strategy not only outperformed more frequently but also did so with statistical significance across the observed periods. While the absolute differences in CAGR are modest, their consistent direction and significance suggest that expiry-based timing may capture a persistent structural advantage. Thus, even over medium-term horizons, aligning SIP execution with F&O expiry may offer a subtle but repeatable edge for disciplined investors.
IV-B3 Five-Year Investment Returns
The five-year investment horizon offers a more stable perspective on SIP performance by dampening short-term volatility and better capturing broader market cycles. Across the four evaluated five-year periods—2005–2009, 2010–2014, 2015–2019, and 2020–2024—both SIP strategies generated consistent growth, with CAGR values typically ranging between 5% and 9%. Full results are presented in Table III and visualized in Figure 3. Importantly, the expiry-based strategy outperformed the first-day strategy in all four periods, with positive CAGR differentials ranging from approximately 0.08 to 0.16 percentage points. This consistent directional advantage lends credibility to the hypothesis that expiry-aligned SIPs benefit from favorable structural or behavioral market effects around the derivatives expiry. The earliest period (2005–2009), covering the post-GFC rebound, showed the largest outperformance (9.04% vs. 8.91%). Though the return gaps in the subsequent periods were more modest, the fact that all differences remained consistently positive suggests a subtle structural edge rather than random noise. The diminishing gaps also reflect how the compounding nature of long-term investing tends to smooth out entry-point differences over time. The (p = 0.0039 value suggest the statistically significant at the 1% level of significance. The five-year results robustly support the notion that expiry-day SIPs may provide a modest but persistent performance edge over traditional first-day investments. While the absolute advantage is small, its consistency—especially during volatile recovery phases—strengthens the argument for expiry-aware SIP execution. This finding affirms that even over longer horizons, intelligent timing can act as a complementary force to disciplined investing.
Figure 3: Year-wise difference in multi-year SIP returns between EXP-SIP and FTD-SIP
IV-B4 Ten- and Twenty-Year Investment Returns
When SIP strategies are evaluated over extended periods of ten and twenty years, the differences between first-day and expiry-day investments become increasingly negligible, underscoring the dominant role of long-term compounding over market entry timing. In the ten-year investment windows—2005–2014 and 2015–2024—the final corpus values and CAGR figures from both strategies converge closely. The first slot shows CAGR values of 6.76% for the first-day strategy and 6.83% for the expiry-based one; the second records 6.99% versus 7.04%, respectively. In both cases, the differences remain within 0.07 percentage points and are statistically insignificant, as confirmed by high p-values ($=0.068>0.05$) in the associated t-tests. These findings emphasize that over a decade-long horizon, the initial variance introduced by timing the SIP installments is gradually averaged out by the overall market trajectory. This convergence becomes even more pronounced in the twenty-year horizon from 2005 to 2024—the longest duration in this analysis.The CAGR difference stands at a minimal 0.03 percentage points (6.65% vs. 6.68%), again insignificant. Over this extended timeline, both strategies effectively absorb all short-term market fluctuations, corrections, and recoveries, leading to nearly identical outcomes.
These results affirm the foundational principle of long-term investing: consistency and discipline outweigh timing precision. While shorter windows may show minor advantages for expiry-based strategies—potentially due to systematic end-of-month market behavior—such advantages diminish to near irrelevance as the investment horizon lengthens. For long-term investors committed to systematic equity exposure, this analysis validates that the SIP execution date need not be optimized. Simply staying invested and allowing the power of compounding to unfold proves to be the most critical factor in wealth creation.
IV-C Consolidated Observations
The comparative analysis of SIPs initiated on the first trading day versus the F&O expiry day of each month reveals a nuanced but consistent pattern. Across nearly all years and investment horizons, the expiry-day strategy tends to yield a marginally higher CAGR than the first-day approach. However, this edge is generally modest, with only a few instances—such as in 2003, 2004, 2005, 2013 and 2020—where the difference is both practically large ($>1.5\%$ ) and statistically significant (p $<0.05$).
Short-term investments (1-year horizon) exhibit the greatest sensitivity to entry timing, likely due to heightened market noise and short-lived momentum or volatility clusters around expiry. For example, in 2003 and 2005, expiry-day SIPs benefited from pronounced dips and recoveries surrounding monthly rollovers, leading to an advantage exceeding 2.5% in CAGR. However, this timing premium proves unstable—vanishing or even reversing in years of market crisis (e.g., 2008, 2020) or extreme bullishness with smooth trends (e.g., 2009). As the investment horizon extends—across 3-year, 5-year, 10-year, and especially 20-year windows—the CAGR difference between the two strategies steadily converges toward zero. This result underscores the principle that the SIP mechanism itself—by distributing entry points across time—acts as a volatility smoother, gradually diluting the impact of any single month’s entry price anomaly. For long-term, goal-oriented investors, this suggests that entry timing becomes largely irrelevant beyond a certain time horizon. What matters more is disciplined, uninterrupted investing, rather than tactical entry attempts. Nonetheless, for momentum-oriented or expiry-aware traders, the small but repeatable edge observed in certain phases suggests that expiry-based SIPs can serve as a useful heuristic—though this should be contextualized within broader asset allocation and risk management frameworks.
In essence, both strategies are statistically and practically equivalent in the long run. However, the expiry-based approach may offer a behaviorally aligned advantage for certain investor types, especially those who follow the market calendar closely. The overarching implication remains clear: SIP works best not because of when one invests in a month, but because one invests every month.
V Discussion
V-A The Myth of 12% SIP Returns
Our results challenge the commonly cited benchmark of 12-15% annualized returns from systematic investment plans (SIPs) in index fund [13, 30, 11]—a figure frequently promoted by financial institutions and investment advisors. As detailed in Table III, the actual pre-tax SIP CAGR over the full sample period (2005–2024) ranges between 6.65% and 6.68%, revealing a significant gap between perception and reality. This divergence arises from three key distortions:
1. Survivorship and Selective Period Bias : Industry narratives often highlight early high-growth years particularly 2003–2007, when equity markets delivered over 29% annualized returns—while downplaying subsequent periods of stagnation, volatility, or crisis. This selective reporting parallels the survivorship and hindsight bias discussed in the literature on mutual fund performance illusion [7, 34]. When the full cycle is considered, the compounding effect of subsequent drawdowns sharply dilutes headline figures.
2. Fee Drag and Expense Ratio Compounding : Even our frictionless CAGR estimates overstate real-world outcomes. Once expense ratios (typically ranging from 1.1% to 1.8% annually) are accounted for, the net return to investors declines further. SEBI’s own disclosures suggest the average post-fee CAGR for equity mutual funds in India over the last two decades is closer to 5.2%, underscoring the cumulative impact of management costs [27].
3. The Investor Behavior Gap (IRG) : Investor-specific timing errors—such as panic redemptions during drawdowns or chasing returns after rallies—contribute an additional drag. This gap between the fund’s time-weighted return and the investor’s dollar-weighted return has been extensively documented by [10], who estimate an average investor return gap of 0.7–1% per year. In the context of SIPs, such behavioral frictions manifest as premature discontinuation or misaligned portfolio switches. Together, these factors explain why the frequently marketed 12% figure is more myth than reality. A realistic expectation—net of fees and behaviorally adjusted—would place long-run SIP returns closer to 5–6%, especially for broad-based index or actively managed equity funds. This has direct implications for financial planning, retirement projections, and public trust in investment advisory practices.
V-B Investor Welfare Considerations
The entry timing penalty—while seemingly marginal—scales into a substantial aggregate impact for the Indian mutual fund ecosystem. With approximately 40 million active SIP accounts [2], even a 0.19% annual CAGR delta implies a cumulative wealth erosion of thousands of crore per annum—exceeding several key line items in the Union Budget. This inefficiency stems from three persistent structural asymmetries dominated by institutional investors, algorithmic traders and HNI. Rectifying disparities requires a shift in regulatory and industry design. We propose three interventions:
  1. 1.
    Mandating SIP timing flexibility during onboarding, with expiry-day options;
  2. 2.
    Requiring monthly expiry flow disclosures by AMCs, analogous to FPI flow reports;
  3. 3.
    Publishing investor dashboards that report real SIP returns adjusted for fees, inflation, and discontinuation risk.
Such reforms would enhance transparency, empower investor agency, and reduce the behavioral and structural drag that currently hampers long-term compounding.
V-C Limitations and Research Frontiers
While this study offers significant insights into the temporal dynamics of SIP investments, it also opens several avenues for further investigation. First, the extent of sectoral heterogeneity remains unclear—do timing effects differ across industries (such as IT vs. FMCG) or market capitalizations (large-cap vs. mid-cap)? Second, the geographic generalizability of our findings warrants exploration. It is unknown whether similar expiry-related timing advantages exist in developed markets with differing microstructure dynamics, such as the U.S. or European Union. Finally, an important practical question remains under active exploration by the author: what is the optimal pattern of incremental monthly SIP contributions (e.g., fixed vs. progressive increase) that maximizes long-term compounded returns?
Addressing these frontiers will likely require richer datasets—such as panel-level investment flows and investor demographics—as well as interdisciplinary approaches combining econometrics, behavioral finance, and simulation modeling. Insights from such research could inform more personalized SIP structures, smarter default options, and regulatory innovations to enhance investor welfare.
VI Conclusion
This paper provides a comprehensive analysis of SIP timing strategies in India’s Nifty 50 index, revealing three core insights with broad theoretical and practical implications. First, we find statistically significant short- to medium-term advantages for SIPs executed on F&O expiry days, which outperform first-day investments by an average annual excess return of 0.58% ($\sigma=1.12\%$, $p<0.05$). This outperformance appears linked to structural liquidity distortions, driven by institutional hedging, month-end rebalancing, and predictable retail flow patterns. Second, our results directly challenge the widely circulated claim of 12% SIP returns. The actual pre-tax CAGR over the 2005–2024 period lies between 6.65% and 6.68%, with the gap explained by survivorship bias ( 3.2 percentage points), compounded expense ratios ( 1.1–1.8%), and a behavioral investor return gap ( 0.7%). These findings expose a critical disconnect between investor expectations and realized outcomes, emphasizing the need for greater transparency in return reporting.
In response, we propose targeted regulatory reforms. SEBI should consider mandating expiry-day SIP options across all mutual fund platforms, implementing real-return dashboards that adjust for inflation and costs, and requiring rolling-period return disclosures to avoid misleading point estimates. For investors, expiry-day SIPs offer a simple, cost-free enhancement for short- to medium-term goals, while conventional strategies remain suitable for long-term wealth creation.
Future research should investigate timing effectiveness across sectors and market caps, explore demographic or behavioral drivers of suboptimal SIP execution, and assess the generalizability of these patterns in other emerging and developed markets. Ultimately, while market microstructure offers fleeting advantages to the informed, sustained long-term participation remains the most reliable and powerful strategy—an enduring truth that transcends any tactical timing edge.

### v2

@@ V-B Deconstructing the ”12% Return” Myth: Biases and Real-World Frictions
    Survivorship and Selective Period Bias: Public narratives often anchor on exceptionally high-growth periods (e.g., 2003–2007) while downplaying subsequent stagnation or crises. This selective reporting parallels the survivorship and hindsight bias discussed in the literature on mutual fund performance illusion [7, 43]. When the full cycle is considered, the compounding effect of subsequent drawdowns sharply dilutes headline figures.
@@
    Compounded Fee Drag: Our frictionless estimates do not account for expense ratios (typically 1.1% to 1.8% annually for equity funds in India). Over decades, these fees create a substantial drag on compounded returns, a factor often overlooked in simplified projections.
  3. 3.
    The Investor Behavior Gap: As documented by [17], investors often underperform the funds they invest in due to behavioral errors like panic selling or performance chasing. This ”return gap” further reduces realized wealth.
Factoring in these realities, a more realistic net, behaviorally-adjusted return expectation for a typical equity SIP would be closer to 5–6%. This has profound implications for financial planning, retirement projections, and the fiduciary responsibility of investment advisors.
V-C Implications for Investor Welfare and Policy
The inefficiencies identified—both the suboptimal default timing of SIPs and the inflated return expectations—have a substantial aggregate impact on investor welfare in India. With approximately 40 million active SIP accounts [2], even marginal performance differences result in a cumulative wealth erosion of thousands of crores annually. Rectifying this requires a shift in both industry practice and regulatory design. We propose three concrete interventions:
  1. 1.
    Mandating that all mutual fund platforms and distributors offer flexible SIP timing options, including expiry-day dates, during investor onboarding.
  2. 2.
    Requiring monthly expiry-related fund flow disclosures by AMCs to increase market transparency, analogous to existing FPI flow reports.
  3. 3.
    Publishing official, standardized investor dashboards that report realistic SIP returns adjusted for fees, taxes, and inflation to better manage investor expectations.
V-D Limitations and Future Research
While this study offers key insights, it is subject to limitations that open avenues for future work. Our analysis is confined to the Nifty 50 large-cap index; it remains unclear if these timing effects generalize across different market caps (mid/small-cap) or sectors. Second, the findings are specific to India’s market structure, and their applicability to developed markets with different derivatives cycles warrants investigation. Finally, future research could explore optimal contribution patterns (e.g., fixed vs. progressively increasing amounts) to further enhance long-term outcomes. Addressing these questions will require richer datasets and interdisciplinary approaches to inform more personalized and effective investment strategies.
VI Conclusion
This paper provides a comprehensive analysis of systematic investment timing in the Indian market, offering two primary contributions with significant implications for investors, advisors, and regulators. First, our research moves beyond the simplistic timing doesn’t matter narrative by empirically documenting a statistically and economically significant advantage for SIPs aligned with derivatives expiry days over short-to-medium-term horizons. Crucially, we establish clear boundary conditions for this effect, demonstrating its systematic decay to irrelevance over multi-decade periods where the mathematical force of compounding and the discipline of continuous participation become the dominant drivers of returns.
Our second major contribution is an empirical correction to the widely marketed ”12% return myth.” We demonstrate that realized, long-term, pre-tax index returns are closer to 7%, and that this figure is further eroded by real-world frictions such as fees, biases, and investor behavior gaps. This finding exposes a critical disconnect between industry marketing and achievable outcomes, highlighting an urgent need for greater transparency.
The central implication of this work is a call for a data-driven evolution in investment practice. For the industry, it supports the design of smarter, timing-aware products and more realistic investor communication. For investors, the message is nuanced but clear: while a simple timing switch can provide a marginal, cost-free enhancement for shorter-term goals, the most critical determinant of long-term wealth creation is not the choice of day, but the unwavering discipline of consistent participation. Ultimately, while market microstructure may offer a fleeting edge, it is the steadfast commitment to staying invested that remains the most powerful and reliable strategy for building wealth.
Appendix A: Monthly SIP Dates and Multi-Year CAGR Comparisons
@@ VI Conclusion
TABLE IV: First trading date and F&O expiry date (in parentheses) for each month
From Year To Year Years CAGR (F) CAGR (E) Difference
@@
TABLE V: Multi-Year SIP Investment Performance: First Day vs F&O Expiry
Appendix B: MATLAB Code Snippet for Statistical Validation
The following MATLAB code summarizes the analysis pipeline used:
% Example: Paired t-test and Cohen’s d
[h, p, ~, stats] = ttest(CAGR_E, CAGR_F);
cohens_d = mean(CAGR_E - CAGR_F) / std(CAGR_E - CAGR_F);
% Wilcoxon Signed-Rank
[p_w, ~, stats_w] = signrank(CAGR_E, CAGR_F);
% Stochastic Dominance
[fE, xE] = ecdf(CAGR_E); [fF, xF] = ecdf(CAGR_F);
All data and MATLAB code used for statistical analysis are available from the corresponding author upon reasonable request.
@@ Acknowledgment
Special appreciation is extended to the organizers and participants of the 7th International Conference on Financial Markets and Corporate Finance (ICFMCF 2025) held at the Vinod Gupta School of Management, IIT Kharagpur. The conference offered critical feedback and expert discussion, which directly helped improve the depth and methodological rigor of this manuscript.